OpenStudy (anonymous):

HELP FOR MEDAL!!! Today’s temperature in Barrow, Alaska, and the high temperatures for the next two days are consecutive integers whose sum is -87. Write an equation and solve it step-by-step using transformations to find the three temperatures.

OpenStudy (anonymous):

wait ill hlp in a min

OpenStudy (anonymous):

@thekakashigaming

OpenStudy (displayerror):

Consecutive integers are like the following: \[1, 2, 3, \ldots\] Notice how each successive integer is one more \(\left(+1\right)\) than the previous. Following this same idea, if I let \(x\) be the first temperature: \[x = \text{Today's temperature}\] What would the expression be for tomorrow's temperature? What would the expression be for the temperature of the day after tomorrow?

OpenStudy (anonymous):

@sammixboo

sammixboo (sammixboo):

Okie dokie, so I will give you the formula to solve this problem n + (n+1) + (n+2) = This is what I use, but not sure if it the formula given in lessons. Give me just a second

sammixboo (sammixboo):

n + (n+1) + (n+2) = -87*

sammixboo (sammixboo):

Where N is your lowest consecutive number and the next two numbers would be the next two higher numbers after that.

OpenStudy (anonymous):

wait how do i find my lowest consecutive number?

OpenStudy (anonymous):

@sammixboo @acxbox22 @Directrix @SolomonZelman